### v2

# ReFRAME or Remain: Unsupervised Lexical Semantic Change Detection with Frame Semantics

Abstract
The majority of contemporary computational methods for lexical semantic change (LSC) detection are based on neural embedding distributional representations. Although these models perform well on LSC benchmarks, their results are often difficult to interpret. We explore an alternative approach that relies solely on frame semantics. We show that this method is effective for detecting semantic change and can even outperform many distributional semantic models. Finally, we present a detailed quantitative and qualitative analysis of its predictions, demonstrating that they are both plausible and highly interpretable.
ReFRAME or Remain: Unsupervised Lexical Semantic Change Detection with Frame Semantics
Bach Phan-Tat1, Kris Heylen1,2, Dirk Geeraerts1, Stefano De Pascale1,3, Dirk Speelman1 1Department of Linguistics, KU Leuven 2Instituut voor de Nederlandse Taal 3Vrije Universiteit Brussel Correspondence: ttbach.phan@kuleuven.be
1 Introduction
Lexical semantic change detection is well-established in Natural Language Processing (NLP), with different shared tasks for different languages, amongst which the most popular one is SemEval 2020 Task 1 Schlechtweg et al. (2020). Most modern systems are based on either static word embeddings like word2vec Mikolov et al. (2013) or ELMo Peters et al. (2018), or contextualised ones such as BERT Devlin et al. (2019) or S-BERT Reimers and Gurevych (2019), the latter currently achieving the state-of-the-art (SOTA) results Cassotti et al. (2023). Despite architectural differences, these methods share a common assumption: the distributional hypothesis Firth (1957) , and can capture some aspects of lexical meaning (e.g., Iacobacci et al., 2016; Lenci et al., 2022; Pilehvar and Camacho-Collados, 2019; Yenicelik et al., 2020). Using them for modelling semantic change would therefore be the logical next step.
Figure 1: Changes in the frequencies of 6 frames that the English noun plane participated in between the 2 subcorpora for English of the SemEval 2020 shared task 1 Schlechtweg et al. (2020). Plane is annotated as semantically changed in the dataset
However, co-occurence is only one perspective of lexical semantics. There are other perspectives/frameworks such as Behavioural Profiles Gries and Divjak (2009), Frame Semantics Fillmore (1982), Conceptual Metaphor Lakoff and Johnson (1981) (see Geeraerts (2010) for a comprehensive overview of lexical semantics), and some of them have also been used in diachronic lexical semantics and semantic change research (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022).
Exploiting Frame Semantics to detecting lexical semantic change is the focus of this paper. Consider the English noun plane: in the 20th century, after the birth of the modern airplane in 1903, the aircraft meaning emerged and became more dominant while the dimensional meaning became less dominant. This is reflected in the shift in the frame distributions of plane, where aircraft-related frames such as commerce_sell, departing, self_motion increase in frequency and dimensional-related frames such as identicality, part_whole decrease in frequency, as shown in Figure 1. Our main hypothesis is that the degree of semantic change of a word is reflected in changes in the frames that word participates in, either as the frame trigger or as a frame element. Note that trying to set a new SOTA is NOT our goal, instead we want to bring more attention to more simple, transparent techniques that utilise explicit linguistic information.
We investigate this question experimentally [^1] All the codes for parsing, preprocessing and experimenting are included in the supplementary materials. using the English dataset [^2] We could only test this method on the English dataset due to the lack of available and reliable frame semantic resources for other languages. This is justified in section 3. from SemEval 2020 Task 1. Our main findings are the following:
  • Tracing the changes in the distribution of frames outperforms many distributional semantic models, including both static embeddings (count-based and prediction-based) and contextualised embeddings.
  • Because the method is interpretable by design, we can decompose change scores into contributions of individual frames and inspect supporting contexts, enabling targeted qualitative analyses to understand how a word has changed (as in the plane example above) in contrast to the opaqueness of dense embedding models [^3] While we acknowledge the growing field of explainable AI and its diverse range of techniques (e.g., Integrated Gradients Sundararajan et al. (2017), LIME Ribeiro et al. (2016) or SHAP Lundberg and Lee (2017), see surveys in Arrieta et al. (2019); Madsen et al. (2022)), these methods are often post-hoc, add methodological and computational overhead, and may not yield explanations that are straightforward to inspect linguistically. This makes our approach particularly convenient for linguistic and humanities research, where direct, feature-level inspection is desired..
  • Our approach is naturally aligned with linguistic theories of lexical semantics and semantic change (e.g., Dalpanagioti, 2019; Geeraerts, 1997, 2010; Geeraerts et al., 2023; Hanks, 2013), helping to bridge NLP methods with diachronic semantics and linguistic analysis.
2 Related work
A long-standing and influential tradition in diachronic semantics follows the Distributional Hypothesis Firth (1957). The rise of corpus linguistics and concordancing makes this idea a practical methodology. Researchers began tracking a target word’s prominent collocates and recurring patterns across corpora and later across time slices, often using association measures for identifying characteristic collocations Church and Hanks (1990); Sinclair (1991). This approach naturally extends to diachrony, where time-sliced collocational profiles reveal how highly associated collocates emerge, fade, or are replaced across periods. The prominence of this line of work is reflected by the many case studies (e.g., Alba-Salas, 2007; Pettersson-Traba, 2021), methods Anonymous_Author (2026); Garcia and García Salido (2019); McEnery et al. (2019) and tools Anthony (2014); Brezina et al. (2020); Kilgarriff et al. (2014); Jurish (2015); Jurish and Nieländer (2020) developed around it. With advances in computational infrastructure and modelling, attention increasingly shifted to a second realisation of the distributional hypothesis, where co-occurrence information is encoded as numerical vectors (embeddings). Early approaches built count-based co-occurrence matrices Bullinaria and Levy (2007); Gulordava and Baroni (2011); Schütze (1998). Later work introduced predictive neural embeddings that compress distributional information into dense vectors, with better performance in capturing semantics Baroni et al. (2014); Bojanowski et al. (2017); Levy and Goldberg (2014); Mikolov et al. (2013); Pennington et al. (2014). More recently, contextualised models extend this paradigm by modelling individual instances of a word Devlin et al. (2019). For semantic change problems, standard workflows split data into time slices, learn word representations per slice, align the spaces, and quantify change by comparing the aligned representations (e.g., via cosine distance or clustering) Tahmasebi et al. (2021); Tahmasebi and Dubossarsky (2023). As LSC research moved toward standardised evaluation with shared tasks and datasets, compact vector representations with a single change score became more attractive. As a result, most work in the area, including the current SOTA, follow this approach (e.g., Cassotti et al., 2023; Geeraerts et al., 2023; Giulianelli et al., 2020; Kim et al., 2014; Kulkarni et al., 2015; Kutuzov et al., 2018; Montariol et al., 2021; Tahmasebi and Dubossarsky, 2023; Tahmasebi et al., 2021).
Other theory-driven frameworks have also been employed in Diachronic Semantics (e.g., Geeraerts, 1983; Jansegers and Gries, 2020; Law, 2019; Pettersson-Traba, 2016; Sovran, 2004; Sullivan, 2007; Sweetser, 1990; Vais, 2022). However, these approaches often yield descriptive outputs, making it difficult to compare systematically across languages and datasets. Recently, efforts have been made to benchmark these theory-driven approaches with standard shared tasks (e.g., Kutuzov et al., 2021; Ryzhova et al., 2021 on grammatical profiles; Tang et al., 2023 on sense distributions). Their strong performances on shared-task benchmarks (sometimes even outperform many embedding-based methods) inspired us to test the effectiveness of using Frame Semantics information in capturing word meaning change.
Frame Semantics was introduced by Fillmore (1982) as a theory of meaning in which words are understood relative to structured background knowledge called frames. A frame is a conceptual scene (e.g., Commercial_transaction, Motion) that specifies the key participants, relations, and typical inferences involved. The meaning of a lexical unit is not just a standalone definition, but a particular way of evoking a frame and profiling parts of it. For example, buy, sell, pay, and cost all relate to the same commercial frame, but each highlights different roles and perspectives. Frames come with frame elements that represent the core components of the scene (e.g., buyer, seller, goods, money). The theory is operationalised in resources such as FrameNet Baker (2014), which catalogs frames, frame elements, and the lexical units that evoke them, along with annotated examples. Because of its structured representation of ‘who did what to whom’ at the level of conceptual situations, Frame Semantic representations have supported applications in semantic parsing, information extraction, and linguistic analysis (e.g., Law, 2019; Li et al., 2019; Marzinotto et al., 2018; Robin et al., 2023; Vais, 2022).
Some practical weaknesses of Frame Semantics are:
  • The substantial time and manual work needed for building corpora.
  • The lack of coverage for different languages as frame inventories are built from particular languages and cultural/institutional settings, although efforts have been made for cross-lingual applicability Boas et al. (2025).
  • The lack of coverage for historical and out-of-domain data as many senses, lexical items, constructions are missing from the inventory Hartmann et al. (2017); Palmer and Sporleder (2010).
These may explain why FrameNet has mainly been used in specific linguistics case studies Blank (1999, 2003); Georgakopoulos (2018); Koch (1999a, 2008, b); Law (2019); Vais (2022) and not in unsupervised computational approaches to LSC detection. As far as we know, our work is the first one to employ Frame Semantics for this task.
3 Data and tasks
Following SemEval 2020 Task 1 on Unsupervised Lexical Semantic Change Detection Schlechtweg et al. (2020), we formulate the problem as either binary classification or graded ranking, corresponding to subtasks 1 and 2. In subtask 1, given a set of target words, the system predicts whether each word gained or lost at least one sense between two time periods. In subtask 2, the system ranks target words by their degrees of semantic change. Annotating semantic change is challenging because it requires assessing meaning differences across many usage instances from each period. A widely used protocol is DURel Schlechtweg et al. (2018), where annotators compare pairs of sentences containing a target word and judge whether the instances are used in the same sense or in different senses. These pairwise judgments are then aggregated into a final change score, either binary or continuous. This procedure was used to create the SemEval 2020 Task 1 dataset, the most widely used one to compared different LSC detection systems. It covers four languages: English, German, Latin, and Swedish, with manual annotations for both subtasks, and provides a diachronic corpus split into two time periods for each language. Using this benchmark enables direct comparison of our approach with SOTA systems under a shared evaluation setup.
Period Tokens Types Corpus Target words
C1 (1810 – 1860) 6.5M Raw, Lemma CCOHA 37
C2 (1960 – 2010) 6.7M Raw, Lemma CCOHA
Table 1: Descriptions of the English dataset in SemEval 2020 Task 1.
However, our approach requires high-quality frame-semantic parses, which are not currently available for all languages. While English benefits from extensive FrameNet resources and several mature automatic frame-semantic parsers, we were unable to find comparably reliable off-the-shelf parsers for the other 3 languages. For English, we used the pretrained Frame Semantic Transformer model Chanin (2023), which is trained on FrameNet 1.7 and is reported to achieve SOTA performance. It is released as a Python library, making it convenient for using and reproducing the results. For German, although there are several FrameNet-like resources, they were all relatively dated Boas (2002); Burchardt et al. (2009, 2006); Erk and Padó (2006). The most recent resource and parser we have found is PFN-DE Bick (2022). However, it is only accessible through a hosted interface that only accepts text files up to 10,000 characters, and processing larger volumes requires contacting the provider for a paid processing agreement 86. This makes it impractical to parse 140 million token historical German corpora Schlechtweg et al. (2020) at scale compared to a locally runnable, openly distributed pipeline, and raises concern about transparency and reproducibility. Moreover, it has only been evaluated on modern Twitter data, making its reliability on historical German uncertain. For Swedish, Svenskt frasnät 2.0 (SweFN) Dataset Språkbanken Text (2024) provides a FrameNet-style resource. The latest attempt to build a pretrained Swedish frame parser with SweFN used the same architecture as the English ones mentioned above, but with substantially lower performance Dannélls and Virk (2025); Dannélls et al. (2024). In a small sanity check with 30 random, diverse and basic sentences, we also observed 13 questionable frame predictions, such as Cutting for Jag såg en film igår‘I saw a movie yesterday’ (see the results in table 4 in the Appendix). The low F1 scores, coupled with its failure in such simple examples, raise concern about the reliability of the model. For Latin, we did not find any public frame-semantic resources or pretrained parsers.
Because of the lack of available and quality frame-semantic parsers for German, Swedish and Latin, we restrict the experiment to English where strong pretrained parsers exist. We compensate for the lack of multilingual quantitative experiments with a detailed qualitative analysis. A quick description of the English data is presented in Table 1. Note that although we evaluate only English, frame semantics is in principle language-agnostic Fillmore (1982) so the approach should generalise once suitable frame resources and parsers exist.
4 Method
4.1 FrameNet parsing
Because we only need sentences containing the target lemma, we first split both the lemmatised and raw data into lemma-specific subcorpora. Since the lemmatised and raw files contain the same number of sentences in the same order, we do this by scanning the lemmatised file to select sentences that include the target lemma, then retrieving the corresponding sentences from the raw (token) file by index. We then frame-parsed them using the Frame transformer model Chanin (2023). The parser is available in two variants, base and small, with similar reported performances. It has also been noted that the models can occasionally identify a frame but return empty outputs. We therefore use base by default and fall back to small when it fails. If both models fail, we skip the sentence. Across 49,416 sentences, we triggered the fallback for 2,413 lemmatised sentences (4.9%) and 1,052 raw sentences (2.1%). Amongst these, both models failed (and the sentences were skipped) for only 805 lemmatised sentences (1.6%) and 336 raw sentences (0.6%).
One technical limitation of the parser is that it returns the frame elements as raw strings without locations. This makes it impossible to locate the target lemma in the parsed token data (one target lemma may appear in different surface forms). We therefore use lemmatised corpora rather than raw corpora. To assess the impact of doing so, we compared (within each target word and time period) the frame distributions derived from the lemmatised and raw corpora using Jensen–Shannon Divergence Menéndez et al. (1997). The resulting JSD values (in the table 5 in the Appendix) are uniformly small (all < 0.15, most < 0.10), indicating that using lemmatised corpora introduces only minor changes to the induced frame distributions.
4.2 Basic procedure
Listing 1: Pseudo-code for how we collect the frames
for sentence in sentences:
for frame in sentence.frames:
if target_lemma == trigger:
append(frame)
continue
for fe in frame.frame_elements:
if target_lemma in fe:
append(frame)
break
To obtain the frame distributions, we collected all the frames in which the target lemma is either the frame trigger or is amongst the frame elements. At the frame instance level, we avoid duplication to prevent inflated counts. Each frame instance is counted once even if 1) the target lemma is both the trigger and a frame element in the same frame, or 2) the target lemma occurs in multiple frame elements of the same frame. At the sentence level, if the target lemma participates in multiple frames within a sentence, all such frames are included. Similarly, multiple (even similar) frame instances within the same sentence are counted as separate occurrences.
We then experiment with 2 types of frame distributions [^4] We intentionally excluded the frame – trigger only distribution. This is because some lemma only appear in frame elements and not evoke any frame, making using frame – triggers only yield sparse or even empty distributions for some targets, causing JSD to be unstable or even undefined.: frame–elements only (FE) and frame–triggers + frame–elements (FTFE). To quantify the distributional changes across periods, we use Jensen–Shannon divergence (JSD) Menéndez et al. (1997). JSD is a symmetric, smoothed variant of Kullback–Leibler Divergence Kullback and Leibler (1951), and with log base 2 it is bounded between 0 and 1, making scores easy to interpret and comparable across targets. In addition, JSD can be decomposed into per-item contributions, allowing fine-grained attribution of which items (in this case, frames) drive the observed change. Prior work on grammatical profiles Kutuzov et al. (2021); Ryzhova et al. (2021) quantified changes with cosine distance between time-specific vectors, but cosine compares vector geometry (angles) rather than directly comparing probability distributions.
5 Results
We evaluate our method on both subtasks of SemEval 2020 Task 1. As described in Section 3, subtask 1 is a binary classification task evaluated by accuracy, while subtask 2 is a ranking task evaluated by Spearman’s rank correlation between the JSDs and the GOLD SCORES. Our primary focus is Subtask 2. For Subtask 1, we classify a word as changed if its JSD >= 0.5, else unchanged.
5.1 Subtask 2
The results for subtask 2 are reported in table 2. Using frame-elements only, we were able to achieve a spearman correlation of 0.249. This is already much higher than the Count Baseline (0.022) and surpassing many static and contextualised embedding systems. Using the combined frame-triggers and frame-elements boosts the score to 0.306, making our system amongst the top 10 systems for English. This is very impressive given the simplicity of the approach.
Team EN score System type
XL-Lexeme 0.757 token
NLPCR 0.436 token
UG_Student_Intern 0.422 type
cs2020 0.375 type
UWB 0.367 type
Discovery_Team 0.361 ens.
Jiaxin & Jinan 0.325 type
UCD 0.307 graph
FrameNet-FTFE 0.306
IMS 0.301 type
Life-Language 0.299 type
Syntactic Dependency 0.277
Entity 0.250 type
FrameNet-FE 0.249
RPI-Trust 0.228 type
Grammatical Profiling 0.218
Random 0.211 type
Skurt 0.209 token
RIJP 0.157 type
UiO-UvA 0.136 token
UoB 0.105 topic
cbk 0.059 token
NLP@IDSIA 0.028 token
Count Bas. 0.022
JCT 0.014 type
TUE -0.155 token
DCC -0.217 type
Freq. Bas. -0.217
Maj. Bas.
Table 2: Performances our models in comparison with other systems in graded change detection (SemEval 2020 Subtask 2), Spearman rank correlation coefficients. FTFE = Frame-Trigger-Frame-Element, FE = Frame Element.
5.2 Subtask 1
The results for subtask 1 are presented in table 3. As the experiment set-up is fully unsupervised, different teams come up with different strategies to classify the changing status of the target lemmas. We based our binary classification score assignment on the 0.5 cutoff point of JSD. The threshold of 0.5 was chosen empirically, based on our prior experience using JSD on linguistic data. It acts as a high-pass filter for semantic salience: it highlights substantial structural shifts while suppressing minor fluctuations that often arise from sparse distributions. We avoid any supervised finetuning for selecting the optimal JSD threshold to preserve the fully unsupervised spirit of the task and comparability with other unsupervised approaches. Both of our systems perform equally well, with an accuracy of 0.622, on par with different word embedding systems.
6 Qualitative analysis
A key strength of our system is its interpretability. JSD decomposition reveals which distributional shifts drive the change, and users can inspect the corresponding contexts to locate the underlying semantic change. In this section, we qualitatively analyse the predictions of our better system (FTFE) to investigate what has been captured.
Team EN score System type
NLPCR 0.730 token
Life-Language 0.703 type
Entity 0.676 type
Jiaxin & Jinan 0.649 type
RPI-Trust 0.649 type
DCC 0.649 type
JCT 0.649 type
Syntactic Dependency 0.649
FrameNet-FTFE 0.622
FrameNet-FE 0.622
UWB 0.622 type
Grammatical Profiling 0.622
NLP@IDSIA 0.622 token
UCD 0.622 graph
Count Bas. 0.595
cs2020 0.595 token
UG_Student_Intern 0.568 type
Skurt 0.568 token
Discovery_Team 0.568 ens.
TUE 0.568 token
Maj. Bas. 0.568
cbk 0.568 token
UoB 0.568 topic
IMS 0.541 type
UiO-UvA 0.541 token
RIJP 0.541 type
Random 0.486 type
Freq. Bas. 0.432
Table 3: Performances of our models in comparison with other systems in binary change detection (SemEval 2020 Subtask 1), accuracy. Note that in this paper we mostly focus on ranking (Subtask 2). FTFE = Frame-Trigger-Frame-Element, FE = Frame-Element.
We follow the same strategy as in our previous study Anonymous_Author (2026), where we divided the target lemmas into 5 groups: True Positive (TP), True Negative (TN), False Positive (FP), False Negative (FN), MID. We define TP as lemmas that fall in the top 33% of both the Subtask 2 Gold Scores and our system’s predicted rankings, and TN as those in the bottom 33% of both. FP are lemmas in the bottom 33% of the gold ranking but the top 33% of our ranking, and FN are lemmas in the top 33% of the gold ranking but the bottom 33% of our ranking. All remaining lemmas are assigned to a MID group. Our qualitative analysis focuses on TP, TN, FP, and FN, as these categories most clearly expose the method’s strengths and limitations. Due to space constraints, we present one illustrative (most extreme) example per group. The lemmas of each group are listed in Table 6 in the Appendix.
Figure 2: The JSD contributions of different frames of the English noun prop. Green bars indicate an increase in relative frequency; red bars indicate a decrease.
True positives
Upon qualitative analysis, semantic changes do indeed link to changes in frame distributions. The English noun prop develops a new theatre object meaning whereas in the earlier period the support meaning (either literal or figurative) was dominant. In the first period, prop were mostly linked to such frames as Possibility (e.g., the possibility of finding support) or Opinion (e.g, thinking about the pillar/bread-winner of the family, or opposing the election of the prop/council of the city) or Building. In the second period, those frames decrease and there was an increase in Self motion frames, where prop acts as the mover (e.g., … props fly into scene) or looking for the prop is the purpose of the mover (e.g., he clambered … looking for perfect props) or Quantified_mass where prop was quantified by few, many or Bringing where prop is the object being brought. This is illustrated in figure 2.
Figure 3: The JSD contributions of different frames of the English noun tree. Green bars indicate an increase in relative frequency; red bars indicate a decrease.
Figure 4: The JSD contributions of different frames of the English noun quilt. Green bars indicate an increase in relative frequency; red bars indicate a decrease.
True negatives
The JSD score of TNs are really low, suggesting that there are minor changes in the frames they appear in. The major shift of the English noun tree (see figure 3) is the fact the phrase old tree appears less (C1: 76 times, C2: 9 times). Other shifts are minor (hence the small overall JSD) and simply contextual (e.g., the decrease in trees being destroyed), not related to any real semantic change.
False positives
We only found one FP (quilt_nn). Our method detects large shifts in its frame distributions, yet it is ranked low in the gold Subtask 2 ranking. This is because of the large shift in the contexts (which do not correspond to any semantic change) in which quilt appears in. In the second period, quilt was the object of many activities, such as lean into the quilt. wring the quilt, lay the folded quilt, etc. This results in a large increase of such frames as Posture, Manipulation, Placing, Spatial contact, Causation, which drives the overall JSD up (see figure 4).
Figure 5: The JSD contributions of different frames of the English noun ball. Green bars indicate an increase in relative frequency; red bars indicate a decrease.
False negative
FNs are target words whose frame JSD scores underestimate the degree of change relative to the gold ranking. For example, our system does capture the shift for the English noun ball from a usage dominated by social event to one dominated by spherical object, reflected in decreases in frames such as Social event and Building subpart (e.g., ball room) and increases in frames such as Cause motion and Manipulation (see figure 5. We hypothesise that, for FN cases, these distributional shifts are just simply not large enough to place the target high in the ranked list. Nevertheless, the method remains useful for qualitative analysis, since it highlights the relevant frame shifts and supports interpretation of semantic change through the lens of Frame Semantics.
There are two common problems we observe during the qualitative analysis. First, some frames are highly general and carry little specific semantic content. For instance, in the ball case, the rise of the Statement frame largely reflects the frequent occurrences of say in contexts that mention ball (whether referring to the social event or the object), rather than a substantive meaning shift. Second, using lemmatised corpora can occasionally disrupt grammatical agreement or sentence boundaries, which may confuse the parser and lead to incorrect frame assignments (this effect appears minor overall, as shown by the low JSD between the raw and lemmatised corpora’s frame inventories).
7 Conclusion
This paper, along with other linguistically motivated case studies on SemEval 2020 Task 1 (e.g., Anonymous_Author, 2026; Kutuzov et al., 2021; Ryzhova et al., 2021; Tang et al., 2023) demonstrates that explicit linguistic evidence is sufficient to build an effective lexical semantic change detection system. Our method achieves competitive performance on English and can outperform many distributional baselines on both subtasks. Although it does not match the current SOTA systems which are based on large pretrained language models and can encode a broader range of linguistic and contextual signals, maximising benchmark scores has never been our primary goal. Rather, we aim to highlight and thus bring more attention to theory-driven, interpretable approaches to lexical semantic change detection.
To further test and refine this approach, future work should:
  • Develop Frame inventories and more reliable Frame parsers for other language
  • Work around the problem of not being able to locate the location of frame elements.
More broadly, we view this approach and some previous ones (e.g., Kutuzov et al., 2021; Anonymous_Author, 2026; Ryzhova et al., 2021) as instances of a general framework for semantic change analysis: identify interpretable, theory-driven dimensions of a target’s distribution (here, frames) and quantify change within each dimension. This framework could (and should) be used as a complementary tool to interpret and validate the changes detected by neural networks.
Limitations
We highlight two primary limitations and practical considerations of our proposed method.
First, frame inventory and frame parsers availability are a hard constraint, making this method inapplicable to languages lacking robust frame parsers. Even for English, the frame parser is still not perfect. It cannot output the token offsets for frame elements, meaning we must use the lemmatised version of the corpora, which can (minorly) affect the performance of the approach.
Second, our qualitative analysis, while supporting the interpretability claim, simultaneously reveals one inherent limitation of distributional frame tracking. The method can overestimate changes when large shifts in a target’s contexts do not correspond to genuine semantic change. Conversely, it can underestimate change when genuine semantic innovations are not sufficiently reflected in frame distributional shifts.
Ethical statement
The data and all codes used in this study are publicly available. No human data was used, and no human participant was involved in the experiment.
Acknowledgments
All ideas and core content were developed by the authors then later polished and optimised using ChatGPT. We reviewed all AI-generated content thoroughly and take full responsibility for its accuracy.
Appendix
Appendix A Testing the Swedish model on 30 sentences
Swedish English translation Predicted verbal frames - SWE Predicted verbal frames - ENG
Jag såg en film igår I saw a movie yesterday Cutting Perception_experience
Vi äger ett hus We own a house Cause_to_fragment Possession
Vi lånar pengar We borrow money Borrowing Getting
Mamma skickar ett brev Mom sends a letter Passing Sending
Barnet dricker mjölk The child drinks milk Ingestion Ingestion
Hon säger sanningen She tells the truth Deciding Telling
Jag vet svaret I know the answer Health_status Awareness
Bilen stannar här The car stops here Halt Halt
Han öppnar fönstret He opens the window Closure Closure
Vi går till parken We go to the park Motion Motion
Tåget anländer nu The train arrives now Arriving Arriving
Boken ligger på hyllan The book is on the shelf Posture No verbal frame. Only Text triggered by book
Jag tänker på dig I think of you Cogitation Awareness
Han älskar sin hund He loves his dog Experiencer_focus Experiencer_focus
Hon förstår frågan She understands the question Frugality Awareness
Vi glömmer namnet We forget the name Memory Remembering_information
De pratar om vädret They talk about the weather Chatting Chatting
Läraren svarar eleven The teacher answers the student Rising_to_a_challenge Education_teaching
Pappa lovar en present Dad promises a present Commitment Cause_to_perceive
Hon ringer sin vän She calls her friend Cause_to_make_noise Referring_by_name
Isen smälter snabbt Ice melts quickly Change_of_phase Change_of_phase
Solen skiner idag The sun shines today Light_movement Light_movement
Han slår igen dörren He slams the door Fall_for Cause_impact
Han skriver ett mejl He writes an email Text_creation Text_creation
Mötet börjar klockan nio The meeting starts at nine o’clock Process_start Process_start
Jag slår på TV:n I turn on the TV Make_noise Change_operational_state
Det regnar ute It is raining outside Precipitation Precipitation
Barnet sover djupt The child sleeps deeply Sleep Sleep
Hon kommer på svaret She comes up with the answer Satisfying Coming_up_with
Hon ger upp She gives up Excreting EMPTY FRAME
Table 4: Testing the Swedish models on 30 random and basic sentences, amongst which 13 frames were questionable (in bold). Compare with the results from the English BASE model..
Appendix B JSD between frame distributions of raw and lemmatised English corpora
Target lemma JSD period 1 (token vs lemma) JSD period 2 (token vs lemma)
attack_nn 0.053278 0.036771
bag_nn 0.093656 0.049945
ball_nn 0.077025 0.076959
bit_nn 0.073138 0.048374
chairman_nn 0.086464 0.042575
circle_vb 0.118216 0.127109
contemplation_nn 0.08533 0.131131
donkey_nn 0.118224 0.097751
edge_nn 0.069459 0.048838
face_nn 0.017471 0.017577
fiction_nn 0.085575 0.065596
gas_nn 0.142413 0.072828
graft_nn 0.126597 0.109724
head_nn 0.017066 0.018624
land_nn 0.02214 0.027298
lane_nn 0.067174 0.065796
lass_nn 0.106964 0.134866
multitude_nn 0.065254 0.130617
ounce_nn 0.122896 0.125564
part_nn 0.03803 0.05156
pin_vb 0.109057 0.102104
plane_nn 0.142745 0.074554
player_nn 0.146468 0.071431
prop_nn 0.134356 0.14812
quilt_nn 0.088169 0.090321
rag_nn 0.09174 0.117761
record_nn 0.068515 0.047556
relationship_nn 0.122012 0.063162
risk_nn 0.118782 0.121072
savage_nn 0.059428 0.128837
stab_nn 0.143965 0.123195
stroke_vb 0.118052 0.089683
thump_nn 0.096097 0.104708
tip_vb 0.092821 0.085674
tree_nn 0.026129 0.035296
twist_nn 0.111025 0.117444
word_nn 0.034585 0.046221
Table 5: JSD (within each target word and time period) between the frame distributions derived from the lemmatised and raw corpora.
Appendix C Predictions of FTFE model
True Positive True Negative False Positive False Negative
prop_nn tree_nn quilt_nn ball_nn
graft_nn face_nn head_nn
stab_nn risk_nn ounce_nn
plane_nn bag_nn
tip_vb
twist_nn
Table 6: True Positive, True Negative, False Positive, False Negative results of FTFE.